The two common chlorides of phosphorus, PCI3 and PCI5, both important in the production of other phosphorus compounds, coexist in equilibrium through the reaction


PCI3(g) + Cl2(g) → PCI5(g)


At 250°C, an equilibrium mixture in a 2.50 L flask contains 0.105g PCI5, 0.220g PCI3 and 2.12g Cl2. What are the values of (a) Kc and (b) Kp for this reaction at 250°C?


Show solutions.


1
Expert's answer
2021-05-04T14:17:34-0400

a) Kc = [PCl5]/[PCl3][Cl2]


with,


molarity = moles/Volume of solution (L)


moles = g/molar mass


we have,


[PCl5] = 0.105/208.24 x 25 = 2.02 x 10^-5 mols


[PCl3] = 0.220/137.333 x 25 = 6.41 x 10^-5 mols


[Cl2] = 2.12/70.906 x 25 = 1.20 x 10^-3 mols


Feed values in Kc equation,


Kc = (2.02 x 10^-5)/(6.41 x 10^-5)(1.20 x 10^-3) = 262.6105


b) Relation between Kc and Kp,


Kp = Kc(RT)^delta(n)


delta(n) = 1 - (1+1) = -1


R = 0.0821 L.atm/K.mol


T = 250 oC = 250 + 273 = 523 K


Feed values,


Kp = 262.6105 (0.0821 x 523)^-1 = 6.116